OpenStudy (anonymous):

The graph shows the system 3x + 6y = 18 and x + 2y = 6. What is the solution for the system?

OpenStudy (callisto):

3x + 6y = 18 3(x+2y) = 3(6) x+2y = 6 therefore, there are infinitely many solutions which satisfies x+2y = 6 ....

OpenStudy (anonymous):

thank you